OpenStudy (anonymous):

2+3=8 3+7=27 4+5=32 5+8=60 6+7=72 7+8=?? solve it ? Brain Teaser

OpenStudy (anonymous):

i will give a medal

OpenStudy (anonymous):

98

OpenStudy (anonymous):

a * (a + b) is the pattern

OpenStudy (anonymous):

Yeah 98

OpenStudy (anonymous):

Sorry it should be a * (a+b) - a

OpenStudy (anonymous):

7 * 15 = 105 - 7 = 98